OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x + 2y = -6 3x + y = 8 A. -2y + 6 B. 2y - 6 C. 2y + 6 D. -2y - 6

OpenStudy (anonymous):

x=2y+6

hero (hero):

-x + 2y = -6 3x + y = 8 2y + 6 = x 3(2y+6)+y = 8

OpenStudy (anonymous):

Thanks :)

hero (hero):

yw :D